# False Discovery Rate and Localizing Power

### v1

@@ Abstract
False discovery rate (FDR) is commonly used for correction for multiple testing in neuroimaging studies. However, when using two-tailed tests, making directional inferences about the results can lead to vastly inflated error rate, even approaching 100% in some cases. This happens because FDR only provides weak control over the error rate, meaning that the proportion of error is guaranteed only globally over all tests, not within subsets, such as among those in only one or another direction. Here we consider and evaluate different strategies for FDR control with two-tailed tests, using both synthetic and real imaging data. Approaches that separate the tests by direction of the hypothesis test, or by the direction of the resulting test statistic, more properly control the directional error rate and preserve FDR benefits, albeit with a doubled risk of errors under complete absence of signal. Strategies that combine tests in both directions, or that use simple two-tailed p-values, can lead to invalid directional conclusions, even if these tests remain globally valid. To enable valid thresholding for directional inference, we suggest that imaging software should allow the possibility that the user sets asymmetrical thresholds for the two sides of the statistical map. While FDR continues to be a valid, powerful procedure for multiple testing correction, care is needed when making directional inferences for two-tailed tests, or more broadly, when making any localized inference.
@@
False discovery rate, two-tailed tests, localizing power, weak control, visualization software.
@@ 1 Introduction
In the context of the multiple testing problem, control over the false discovery rate (fdr) provides an alternative to the control over the familywise error rate (fwer): while fwer refers to the chance of even one type i (false positive) error is committed in a set (family) of statistical tests, fdr refers to the proportion of such errors among tests for which the null hypothesis has been rejected, that is, the proportion of false discoveries among all discoveries. These topics have been extensively covered in the literature; for an early review of the multiple testing problem in brain imaging, see Nichols and Hayasaka (2003); for an overview of the development of fdr, see Benjamini and Hochberg (2000); for its introduction to neuroimaging, see Genovese et al. (2002); for a critique, see Chumbley and Friston (2009); for a discussion of various fdr approaches, see Korthauer et al. (2019). Methods for fwer correction in brain imaging are typically based on the random field theory (rftWorsley et al., 1996, 2004), permutation tests (Holmes et al., 1996; Nichols and Holmes, 2002; Winkler et al., 2014), Monte Carlo simulations (Poline and Mazoyer, 1993; Forman et al., 1995; Cox, 1996) and occasionally, the method by Bonferroni (1936).
Two hallmarks of most neuroimaging studies are that they can involve thousands of statistical tests, and that one can reasonably expect effects to be observed on both directions for many of these. Controlling the fdr aims to provide more power than fwer when many tests are expected to show true effects. This benefit has seen fdr widely adopted in neuroimaging. However, the behavior of fdr when signal manifests in positive direction for some tests, and negative for others, has received little attention. We demonstrate that merely applying fdr does not provide optimal sensitivity when the researcher wishes to make inferences about the direction after using two-tailed tests, and can lead to invalid conclusions in cases of effects asymmetrically distributed throughout the tests.
@@
In this note we revisit the use of fdr in brain imaging, discuss interpretation for two-tailed tests in the light of statistical properties of fdr-controlling procedures, and propose methods that can be considered to address issues that emerge when fdr is used with two-tailed tests. We also bring attention to the fdr approach proposed by Benjamini et al. (2006), which is more powerful than the one originally proposed by Benjamini and Hochberg (1995), and which has seldom been used in brain imaging. We also offer a method for p-value adjustment using this more powerful procedure.
@@ 2 Theory
Consider a number $V$ of voxels (or vertices, or edges, or faces) of an image representation of the brain. For each, a null hypotheses $\mathcal{H}_{1}$, $\ldots$, $\mathcal{H}_{V}$ is tested, and a corresponding p-value $p_{1}$, $\ldots$, $p_{V}$ is computed. Each p-value can be considered “significant” if equal or below some pre-defined test level $\alpha$ (typically, $\alpha$ $=$ $0.05$). Simply comparing each p-value to $\alpha$ leads to many of them to be spuriously considered significant, even if all null hypotheses are true; for example, if $\alpha$ $=$ $0.05$, 5% of the tests will are expected to be found as significant even if no effect is present in any of them.
@@
Regardless of the degree of dependency among tests, it may be the case that, faced with the task of screening many such tests, a researcher may be willing to tolerate a small amount of them to be incorrectly declared significant (false discoveries), as long as a great majority of tests are correctly declared as such (true discoveries), and the errors are limited. In other words, the researcher may be willing to accept some false discoveries as long as their proportion over all discoveries does not spill over some pre-defined threshold $q$ (often $q$ $=$ $0.05$). Controlling this proportion is the essence of fdr approaches.
@@ 2.1 The basic fdr procedure: bh
Control over the fdr can be achieved with the procedure introduced by Benjamini and Hochberg (1995) (bh): let $p_{(1)}\leqslant\ldots\leqslant p_{(V)}$ be their ordered p-values. Find the largest index $i$ for which:
@@
where $q$ is the upper limit on the proportion of false discoveries that the researcher is willing to tolerate. Then reject all null hypotheses that have p-values smaller than or equal to that critical $p_{(i)}$. Graphically, one would lay the ordered p-values in a Cartesian plane that has as horizontal axis the index $i$; draw a line with slope $q/V$ passing through the origin, then reject all tests whose ordered p-values lie to the left of the highest p-value at or below that line (Figure 1, upper-left panel).
Figure 1: Comparison of fdr correction methods, for $q=0.20$. Each symbol $\boldsymbol{\times}$ represents an uncorrected p-value; these are placed in ascending order ($i$); the p-values considered significant according to the respective method are surrounded by a circle $\boldsymbol{\otimes}$. In bh, p-values smaller than or equal to the highest p-value below a single inclined threshold (blue line) are declared significant. In bky, the thresholds themselves are adaptive: their inclination increases for larger p-values (colored lines), making it easier to detect significant p-values, provided that smaller ones are also found significant (for clarity, we show in dashed lines the thresholds for tests considered not significant). In by, the bh threshold line is scaled down by a constant factor, becoming less inclined (more conservative); the same principle used in by can be applied to bky.
The p-values used in this example are: {0.0026, 0.01, 0.014, 0.025, 0.042, 0.066, 0.1, 0.12, 0.17, 0.28, 0.36, 0.524, 0.61, 0.68, 0.78, 0.9, 0.96}.
When the tests are independent, the bh procedure controls the fdr at the level $qV_{0}/V$, where $V_{0}$ is the (unknown) number of true null hypotheses. However, complete independence is not strictly necessary: it is sufficient for tests to exhibit a form of dependence that has been termed positive regression dependency on subsets (prds). prds means that, for a set of tests that are positively dependent, the dependence is the same or stronger within arbitrarily defined subsets where the null hypothesis is true. Under this condition, the bh procedure still controls the fdr (Benjamini and Yekutieli, 2001). Typical imaging analyses in which the value of a test does not depend on whether another test is included or not in the analysis satisfy the prds condition.333prds is a generalization of the concept of positive regression dependency (prd) whereby, for two variables, if one in increases, the expected value of the other does not decrease (Sarkar, 1969); for a review of various related measures of dependence, see Kimeldorf and Sampson (1989). In prds, if the set of p-values is partitioned, prd remains satisfied or is strengthened within each partition.
2.2 A more robust fdr procedure: by
A procedure that controls the fdr more generally, for any arbitrary dependence between tests (as opposed to only under independence or prds) was proposed by Benjamini and Yekutieli (2001) (by). The procedure is similar to bh, except from a small modification in Equation 1, which becomes $p_{(i)}\leqslant iq/V/c(V)$, where $c(V)=\sum^{V}_{i=1}1/i$. Graphically, this results in a less inclined threshold line, and thus, a smaller cutoff $p_{(i)}$ (more conservative) than with the original bh procedure (Figure 1, lower-left panel).
@@ 2.3 A more powerful fdr procedure: bky
Unfortunately, the above procedures are both conservative in the sense that the fdr is, on average, below $qV_{0}/V$, which is always less than or equal to the desired level $q$. To mitigate this conservativeness, Benjamini et al. (2006) proposed to estimate $V_{0}$ after a first pass using bh with $q^{\prime}=q/(1+q)$, then repeat with $q^{*}=q^{\prime}V/\hat{V}_{0}$ if at least one hypothesis was rejected in the first pass; this two-pass procedure has been proposed for use in brain imaging (Chen et al., 2009), albeit it remains rarely used. In their original paper, Benjamini et al. (2006) recognised that there is no reason to stop after the second pass; in fact, the procedure can be repeated iteratively until there are no more rejections. Doing so leads to following procedure (bky), which we adopt here: starting from the lowest p-value, keep rejecting hypotheses for the successive $p_{(i)}$ as long as at least one $p_{(j)}$, $j\geqslant i$, satisfies:
@@
The bky procedure was developed as an improvement over bh, and thus the same considerations hold regarding dependencies between tests. The procedure can also be formulated taking by as starting point; the result is that the bky threshold is scaled by $1/c(V)$.
As with bh and by, bky can also be interpreted graphically (Figure 1, upper-right panel): for each ordered p-value $p_{(i)}$, trace a threshold line with slope $q/(V+1-i(1-q))$ passing through the origin, and check if any p-value from that position (inclusive) onward is at or below that line; if so, reject the corresponding null hypothesis, and move to the next p-value, now with a new threshold line that has a steeper (more liberal) threshold than the previous. The bky procedure differs from bh, however, not only on the use of an adaptive threshold line, whose slope varies according to the position $i$, but also on the stopping criterion: whereas bh uses as cutoff the largest $p_{(i)}$ that satisfies its inequality, bky stops at the first $p_{(i)}$ that no longer satisfies it. This has importance when computing adjusted p-values, discussed next.
@@ 2.4 Adjustment of p-values
It is convenient to inspect a map of p-values that incorporates correction for multiple testing: thresholding such map at the desired test level immediately reveals significant results and is straightforward to examine in most imaging software via, e.g., a slider, or by varying the upper or lower limits of the color scale, or by changing image contrast window center and width. In contradistinction, maps of uncorrected p-values need be thresholded at some corrected $\alpha$ or $q$ levels. Corrected p-values for the bh procedure can be found by solving Equation 1 for $q$ at every position $i$:
@@
where the inequality becomes equality at the lowest (strictest) value of $q_{(i)}$ for which the respective $p_{(i)}$ is significant. Likewise, for bky, corrected p-values can be found by solving Equation 2 for $q$ at $i$:
@@
Unfortunately, corrected p-values for bh and bky are not guaranteed to be monotonically related to their corresponding uncorrected p-values. This can be visualized as occurring due to the fact that they correspond, graphically, to the slope of the line that connects each point to the center of the Cartesian grid, and that slope can be higher than $q$ even for p-values considered significant. Such lack of monotonicity is problematic because thresholding a map of such corrected p-values at some desired level $q$ will not produce the same result as thresholding the map of original p-values at the critical $p_{(i)}$ identified with the respective fdr-controlling method. Following Yekutieli and Benjamini (1999, Definition 2.4), adjusted p-values, which replace corrected p-values without having this issue, can be computed for the bh procedure as:
@@
Figure 2: Comparison of fdr-adjusted p-values produced with different methods, for the same data shown in Figure 1. Each symbol $\boldsymbol{\times}$ represents a uncorrected p-value; these are placed in ascending order ($i$); the p-values considered significant are surrounded by a circle $\boldsymbol{\otimes}$. The adjustment ensures that a single, fixed threshold can be applied to all tests (blue line; $q=0.20$); values below that threshold are considered significant. Note that adjusted p-values can sometimes be greater than 1; these can be replaced by 1 (markers $\boldsymbol{\times}$ in red).
With adjusted p-values, significance thresholds no longer depend on the position $i$ for bky: maps of adjusted p-values can be arbitrarily thresholded at any $q$ in the interval between 0 and 1 (Figure 2). Using Equation 6 for bky, as opposed to a formulation equivalent to Equation 5, solves the issue with the paradoxical results observed by Reiss et al. (2012), whereby adjusted p-values would seem smaller than their unadjusted counterparts.
@@
Note that adjusted p-values computed for various fdr procedures or for Bonferroni can be greater than 1; this can be treated by replacing adjusted p-values greater than 1 by 1 (i.e., by Winsorization of adjusted p-values at 1; in Figure 2 these Winsorized, adjusted p-values are marked in red).
2.5 Strong and weak control
Methods for controlling fwer and fdr differ not only on the measurement of error that is controlled, but on at least another aspect: the methods used for fwer in imaging provide strong control over the error rate, whereas those for fdr provide weak control. Strong control means that the error rate is controlled for any configuration of true and false null hypotheses; in contrast, weak control pertains to controlling the error rate under certain conditions, such as the complete null (Hochberg and Tamhane, 1987; Nichols and Hayasaka, 2003; Proschan and Brittain, 2020). With fdr, weak control ensures that the average proportion of erroneous rejections, given that any rejections occur, is limited to a pre-specified level. While strong control guarantees a bound on errors irrespective of the underlying truth, weak control emphasizes average performance, which can be more lenient but also adaptive to the amount of effects present in the data. In the case of fdr, this has important yet little appreciated consequences on how statistical maps are visualized and interpreted, particularly when using two-tailed tests, as discussed below.
2.6 Discrete p-values
Permutation and other resampling-based methods for inference produce p-values that are discrete. In a permutation test, for example, if a number $J$ of permutations is performed, the smallest possible p-value is $1/J$; all other p-values are multiples $1/J$. Further, if multiple tests are performed, a large number of ties (identical p-values) are expected to occur. The fdr-controlling procedures discussed here make use of the rank $(i)$ of p-values. Even though in many rank-based statistical tests, ties are replaced by their average rank (Hollander et al., 2014), this is not necessary with these procedures: tied p-values can be sorted randomly within their rank; if any p-value is rejected, regardless of its rank, all others tied up with it will also be rejected. While ties are not of concern, the discreteness causes these procedures to become conservative; the degree of conservativeness depends on the granularity of p-values: the less resolved p-values, the more conservative the correction becomes. The reason is that is may be impossible for some p-values reach the threshold at position $i$ simply because they cannot be small enough. This problem can be solved by running more permutations if the data and experimental design permits. Addressing the fdr conservativeness when the number of permutations is very small and cannot be increased remains an unsolved problem. Of note, fdr-controlling procedures for cases in which the different tests may have p-values with varying degrees of resolution have been proposed (for a review, see Döhler et al., 2018); these do not typically apply to imaging, since under the null, the resolution of the permutation distribution of p-values is the same for all tests (a notable exception is if inference uses the negative binomial approximation; see Winkler et al., 2016).
2.7 Two-tailed tests
If the distribution of the test statistic is symmetric and continuous, two-tailed p-values can be obtained from one-tailed as $p^{\text{two}}=1-\left|2p^{\text{one}}-1\right|$. More generally, if the distribution is either continuous or discrete, two-tailed p-values can be obtained as $p^{\text{two}}=2\cdot\min\left\{p^{\text{one}},1-p^{\text{one}}+C\right\}$, where $C=0$ for continuous (often parametric) p-values, or $C=1/J$ for discrete, permutation-based (non-parametric) p-values, and $J$ is the number of permutations. If the distribution of the test statistics is asymmetric, two-tailed p-values must be obtained from that distribution, even if one-tailed p-values are available.
In principle, using fdr with two-tailed tests is straightforward: the two-tailed p-values are obtained, then an fdr-controlling procedure is applied. For visualization, a number of options can be considered: for example, a map of uncorrected p-values can be thresholded at the fdr-controlling level, or a map of adjusted p-values can be thresholded at the desired level. Alternatively, if the distribution of the test statistic is known, the threshold p-value can be converted to a threshold statistic (using the cumulative distribution function), which can be applied to the partition statistical map into significant and non-significant regions. A thresholded map (of statistics or of p-values) can be binarized and used as a mask to display or select regions for further analyses. As long as inference remains focused equally on all tests, using fdr on two-tailed p-values has nothing special compared to using one-tailed p-values. In some circumstances, however, two-tailed tests can become problematic, as discussed below.
2.8 Direction and localization errors

### v2

@@ Abstract
False discovery rate (FDR) is commonly used for correction for multiple testing in neuroimaging studies. However, when using two-tailed tests, making directional inferences about the results can lead to vastly inflated error rate, even approaching 100% in some cases. This happens because FDR control the error rate only globally, over all tests, not within subsets, such as among those in only one or another direction. Here we consider and evaluate different strategies for FDR control with two-tailed tests, using both synthetic and real imaging data. Approaches that separate the tests by direction of the hypothesis test, or by the direction of the resulting test statistic, more properly control the directional error rate and preserve FDR benefits, albeit with a doubled risk of errors under complete absence of signal. Strategies that combine tests in both directions, or that use simple two-tailed p-values, can lead to invalid directional conclusions, even if these tests remain globally valid. To enable valid thresholding for directional inference, we suggest that imaging software should allow the possibility that the user sets asymmetrical thresholds for the two sides of the statistical map. While FDR continues to be a valid, powerful procedure for multiple testing correction, care is needed when making directional inferences for two-tailed tests, or more broadly, when making any localized inference.
@@
False discovery rate, two-tailed tests, localizing power, subsetting property
@@ 1 Introduction
In the context of the multiple testing problem, control over the false discovery rate (fdr) provides an alternative to the control over the familywise error rate (fwer): while fwer refers to the chance of even one type i (false positive) error is committed in a set (family) of statistical tests, fdr refers to the proportion of such errors among tests for which the null hypothesis has been rejected, that is, the proportion of false discoveries among all discoveries. These topics have been extensively covered in the literature; for an early review of the multiple testing problem in brain imaging, see Nichols and Hayasaka (2003); for a review in genetics, also broadly applicable to imaging, see Goeman and Solari (2014); for early papers and for an overview of the development of fdr, see Seeger (1968); Sorić (1989); Benjamini and Hochberg (2000); for its introduction to neuroimaging, see Genovese et al. (2002); for a critique, see Chumbley and Friston (2009); for a discussion of various fdr approaches, see Korthauer et al. (2019). Methods for fwer correction in brain imaging are often based on the random field theory (rftWorsley et al., 1996, 2004), permutation tests (Holmes et al., 1996; Nichols and Holmes, 2002; Winkler et al., 2014), Monte Carlo simulations (Poline and Mazoyer, 1993; Forman et al., 1995; Cox, 1996) and occasionally, the method based on Bonferroni (1936).
Two hallmarks of most neuroimaging studies are that they can involve thousands of statistical tests, and that one can reasonably expect effects to be observed in both directions for many of these. Controlling the fdr aims to provide more power than fwer in the setting where many null hypotheses are false. This benefit has seen fdr widely adopted in neuroimaging. However, the behavior of fdr when signal manifests in positive direction for some tests, and negative for others, has not received as much attention, even if it is a pertinent problem in the field. In principle, using fdr with two-tailed tests is straightforward: the two-tailed p-values are obtained, then an fdr-controlling procedure is applied. As long as inference remains focused on all tests, using fdr on two-tailed p-values has nothing special compared to using one-tailed p-values. Using two-tailed tests can become problematic if, after doing the test, the researcher proceeds to draw inferences on each direction: fdr is not controlled within one direction or another, only globally. Previous literature on directional hypotheses pertinent to this problem can be found in Shaffer (1995, 2002); Williams et al. (1999); Benjamini and Yekutieli (2005). The same problem emerges if a researcher attempts to draw inferences within regions of a map of p-values after applying fdr globally across all of them. Both problems are a general manifestation of the fact that fdr, while a powerful procedure, does not have the ability to allow localization within subsets of tests, that is, it lacks localizing power. As one of our reviewers pointed out, “making discoveries in either direction while controlling the fdr is trickier than it may seem”.
@@
In this note we start by briefly revisiting the multiple testing problem in imaging, p-value adjustment, the use of fdr, and discuss interpretation for two-tailed tests in the light of statistical properties of fdr-controlling procedures. We demonstrate that merely applying fdr does not control the error rate on each side of the statistical map when the researcher wishes to make inferences about the direction after using two-tailed tests, and can lead to invalid conclusions when effects are asymmetrically distributed across the two sides. We investigate methods that can be considered to address such issues, with focus on a slight modification of the method proposed by Benjamini and Yekutieli (2005). We also bring attention to the fdr approach proposed by Benjamini et al. (2006), which is more powerful than the one originally proposed by Benjamini and Hochberg (1995), and which has seldom been used in brain imaging. We also offer a method for p-value adjustment using this more powerful procedure. While the paper is presented in a neuroimaging context, the same topics are relevant to other fields in which fdr is widely used.
@@ 2 Theory
2.1 Preliminaries
Consider a number $V$ of voxels (or vertices, or edges, or faces) of a representation (image) of the brain. For each voxel, a null hypothesis $\mathcal{H}^{0}_{i}$, $i\in\{i,\ldots,V\}$ is tested through the computation of some test statistic $t_{i}$, and a corresponding (one-tailed) p-value $p_{i}$ that represents the probability of finding $t_{i}$ at least as large as the one observed. Each p-value can be considered statistically significant if equal or below some pre-defined test level $\alpha$ (typically 0.05). If the distribution of the test statistic is symmetric around zero and continuous, two-tailed p-values can be obtained from one-tailed as $p^{\text{two}}=1-\left|2p^{\text{one}}-1\right|$. More generally, if the distribution is either continuous or discrete, two-tailed p-values can be obtained as $p^{\text{two}}=2\cdot\min\left\{p^{\text{one}},1-p^{\text{one}}+C\right\}$, where $C=0$ for continuous (often parametric) p-values, or $C=1/J$ for discrete, permutation-based (non-parametric) p-values, and $J$ is the number of permutations. If the distribution of the test statistics is asymmetric, two-tailed p-values must be obtained from that distribution, even if one-tailed p-values are available.
2.2 The multiple testing problem
@@
Regardless of the degree of dependency among tests, it may be the case that, faced with the task of screening many such tests, a researcher may be willing to tolerate a small amount of them to be incorrectly declared significant (false discoveries), as long as a great majority of tests are correctly declared as such (true discoveries), and the errors are limited; if no discoveries are made, there are no false positives by definition. In other words, the researcher may be willing to accept some false discoveries as long as their proportion over all discoveries does not spill over some pre-defined threshold $q$ (often $q$ == $0.05$). Controlling this proportion is the essence of fdr approaches.
@@ 2.3 The basic fdr procedure: bh
Control over the fdr can be achieved with the procedure by Benjamini and Hochberg (1995) (bh): let $p_{(1)}\leqslant\ldots\leqslant p_{(V)}$ be the ordered p-values. Find the largest index $i$ for which:
@@
where $q$ is the upper limit on the proportion of false discoveries that the researcher is willing to tolerate. Then reject all null hypotheses that have p-values smaller than or equal to that critical $p_{(i)}$. Graphically, one would lay the ordered p-values in a Cartesian plane that has as horizontal axis the index $i$; draw a line with slope $q/V$ passing through the origin, then reject all tests whose ordered p-values lie to the left of the highest p-value at or below that line (Figure 1, left panel).
[…]
When the tests are independent, the bh procedure controls the fdr at the level $qV_{0}/V$, where $V_{0}$ is the (unknown) number of true null hypotheses. However, complete independence is not strictly necessary: Benjamini and Yekutieli (2001) showed that it is sufficient that tests exhibit a form of dependence that they termed positive regression dependency on subsets (prds). In this form of dependence, knowledge that one p-value where the null is true is small does not make less likely that the other (V1)(V-1) p-values are also small. If this condition is satisfied, the bh procedure still controls the fdr. Analyses where a false positive does not make the other tests less likely to be significant, as typical in brain imaging, satisfy the prds condition.222prds itself is defined in Benjamini and Yekutieli (2001); additional literature on forms of dependency between variables, relevant to prds, can be found in Lehmann (1966); Esary et al. (1967); Sarkar (1969); Kimeldorf and Sampson (1989).
A procedure that controls the fdr more generally, for any dependence between tests (as opposed to only under independence or prds) was proposed by Benjamini and Yekutieli (2001) (by). The procedure is similar to bh, except for a small modification in Equation 1, which becomes $p_{(i)}\leqslant iq/V/c(V)$, where $c(V)=\sum^{V}_{j=1}1/j$. Graphically, this results in a less inclined threshold line, and thus, a smaller cutoff $p_{(i)}$ (more conservative) than with the original bh procedure.
@@ 2.4 A more powerful fdr procedure: bky
Unfortunately, bh and by are both conservative in the sense that the expected fdr is below $qV_{0}/V$, which is always less than or equal to the desired level $q$. To mitigate this conservativeness, Benjamini et al. (2006) proposed to estimate $V_{0}$ after a first pass using bh with $q^{\prime}=q/(1+q)$, then repeat with $q^{*}=q^{\prime}V/\hat{V}_{0}$ if at least one hypothesis was rejected in the first pass; this two-pass procedure has been proposed for use in brain imaging (Chen et al., 2009), albeit it remains rarely used. In their original paper, Benjamini et al. (2006) recognised that there is no reason to stop after the second pass; in fact, the procedure can be repeated iteratively until there are no more rejections. Doing so leads to following procedure (bky), which we adopt here: starting from the lowest p-value, keep rejecting hypotheses for the successive $p_{(i)}$ as long as at least one $p_{(j)}$, $j\geqslant i$, satisfies:
@@
As with bh, bky can also be interpreted graphically (Figure 1, right panel): for each ordered p-value $p_{(i)}$, trace a threshold line with slope q/(V+1i(1q))q/(V+1-i(1-q)) passing through the origin, and check if any p-value from that position (inclusive) onward is at or below that line; if so, reject the corresponding null hypothesis, and move to the next p-value, now with a new threshold line that is steeper (more liberal) than the previous. The bky procedure differs from bh, however, not only on the use of an adaptive threshold line, whose slope varies according to the position $i$, but also on the stopping criterion: whereas bh uses as cutoff the largest $p_{(i)}$ that satisfies its inequality, bky stops at the first $p_{(i)}$ that no longer satisfies it. This has importance when computing adjusted p-values, discussed next.
@@ 2.5 Adjustment of p-values
It is convenient to inspect a map of p-values that incorporates correction for multiple testing: thresholding such map at the desired test level immediately reveals significant results and is straightforward to examine in most imaging software via, e.g., a slider, or by varying the upper or lower limits of the color scale, or by changing image contrast window center and width. In contradistinction, maps of uncorrected p-values need be thresholded at some corrected $\alpha$ or $q$ levels. Corrected p-values for the bh procedure can be computed by finding the smallest $q$ over every position $i$ that still satisfies the inequality given in Equation 1:
@@
where the inequality becomes equality at the lowest (strictest) value of $q_{(i)}$ for which the respective $p_{(i)}$ is significant (the notation with brackets $[$ ]] avoids double subscripts). In a similar manner, for bky, corrected p-values can be found by solving Equation 2 for $q$ at $i$:
@@
Unfortunately, corrected p-values for bh and bky are not guaranteed to be monotonically related to their corresponding uncorrected p-values (that is, corrected p-values not necessarily increase or decrease as their corresponding uncorrected p-values increase or decrease). Graphically, the corrected p-values correspond to the slope of the line that connects each point to the center of the Cartesian grid, and that slope can be higher than $q$ even for p-values considered significant. Such lack of monotonicity is problematic because thresholding a map of such corrected p-values at some desired level $q$ will not produce the same result as thresholding the map of original p-values at the critical $p_{(i)}$ identified with the respective fdr-controlling method. One may expect monotonicity when doing statistical tests: it allows one to trade-off between maximizing power while minimizing false positives (Zeisel et al., 2011). Following Yekutieli and Benjamini (1999, Definition 2.4), adjusted p-values, which replace corrected p-values without having this issue, can be computed for the bh procedure as:
@@
With adjusted p-values, significance thresholds no longer depend on the position $i$ for bky: maps of adjusted p-values can be arbitrarily thresholded at any $q$ in the interval between 0 and 1 (Figure 2). It should be noted that ties are very common with adjusted p-values; they pose no problem for interpretation. Interestingly, using Equation 6 for bky, as opposed to a formulation equivalent to Equation 5, solves the issue with the paradoxical results observed by Reiss et al. (2012), whereby adjusted p-values would seem smaller than their unadjusted counterparts.
[…]
2.6 Localizing power
Methods for fwer and fdr control differ not only on the measurement of error that is controlled, but on at least another aspect: if a procedure controls the fdr at some level for a set of hypotheses, this control does not necessarily extend to every subset of hypotheses, and the actual fdr can in fact be much higher (or even much lower) for any subset; a procedure that controls the fwer for a set of hypotheses, however, also controls the fwer for every subset. This feature, that fwer procedures have, but fdr procedures have not, was described by Finner and Roters (2001), and called subsetting property by Goeman and Solari (2014). In the context of imaging, we find the term localizing power (or the lack of it) preferable as it immediately conveys the meaning that further localization within parts of an image may or may not be possible, depending on the procedures used for inference (Friston et al., 1996; Nichols and Hayasaka, 2003; Noble et al., 2022).